OpenStudy (anonymous):

Find the x- and y- intercepts of the line that passes through the given points. Then I get this, (4,-2), (5,-4) I don't know what to do with this, can someone please help me?

OpenStudy (hoblos):

do you know how to find the equation of the line passing through these points

OpenStudy (anonymous):

No, thats what i need help with.I don't know how to figure this out at all I need help.

OpenStudy (hoblos):

first get the slope m=(-4+2)/(5-4) = -2 so y+2=-2(x-4) y+2 = -2x +8 2x+y = 6

OpenStudy (hoblos):

for x=0 ---> y=6 (0,6) y-intercept for y=0 ---> 2x=6 x=3 (3,0) x-intercept

OpenStudy (anonymous):

o ok thanks

OpenStudy (hoblos):

my pleasure XD